# End-to-end Contrastive Language-Speech Pretraining Model For Long-form Spoken Question Answering

### v1

@@ Introduction
We use 220M Paraformer (Gao et al. 2022) and BGE-base (Chen et al. 2024) to build CLSR. BGE is frozen during joint training. We consider two models as baseline: one is the E2E text-speech contrastive model like Figure 2, and the other is the cascaded model that first uses ASR model to convert speech into text, followed by a text QA task. For the former, we choose CLAP and SpeechDPR for comparison. For the latter, we use Whisper (Radford et al. 2023) as the ASR module and BGE-base as the text QA module. The Whisper’s size is 244M. In the experiment, word error rate (WER) is used to measure the ASR performance, and top-k question-to-context and context-to-question retrieval recall are used to assess retrieval performance. We establish the experimental environment based on Funasr (Gao et al. 2023) and ModelScope. The $\alpha$ and $\beta$ of the loss is set to $\frac{1}{3}$. We train the model until convergence, consistently using the Adam optimizer with a learning rate of 5e-5.

### v2

@@ Introduction
CLSR is built with Paraformer (Gao et al. 2022) (220M) and frozen BGE-base (Chen et al. 2024) (109M). Two kinds of baselines are compared: an E2E text‑speech contrastive model like Figure 2 and a cascaded model that first transcribes speech with an ASR module followed by a text QA module. For the former, CLAP and SpeechDPR are selected; for the latter, we employ Whisper (Radford et al. 2023) (244M) for ASR and BGE‑base for text QA. Evaluation uses word error rate (WER) for ASR performance and top‑k retrieval recall for retrieval ability. Experiments are conducted on FunASR (Gao et al. 2023) and ModelScope. The loss weights $\alpha$ and $\beta$ are set to $\frac{1}{3}$. Models are trained to convergence with the Adam optimizer at a learning rate of 5e-5.